OpenStudy (anonymous):

Find values of a and b such that f is differentiable everywhere. f(x) = {2cosx if x < pi/2} {ax - b if x > & = pi/2}

OpenStudy (dumbcow):

to be differentiable at pi/2 then it should be continuous and the lim f(x) at pi/2 should exist 2cos(pi/2) = a(pi/2)-b 0 = a(pi/2) -b b = a(pi/2) set the derivatives equal at pi/2 -2sin(pi/2) = a -2 = a --> b = -2(pi/2) = -pi